# Causal Attention with Lookahead Keys

### v1

@@ 2.4 Efficient Inference with UQ-KV Cache
In this section, we introduce the inference algorithm. We first introduce the decoding algorithm. The decoding algorithm consists of the following 2 steps: updating step and combining step. Fix $t\in\left\{1,2,\dots,L\right\}$, and consider generating the $(t+1)$-th token.
@@
The proof of (10) is given in Appendix 8. Next, we discuss the caching strategy and FLOPs of the updating step.
@@
For the prefilling stage, let the prompt length be $L$ and inputs $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$. We first compute $\boldsymbol{\mathit{Q}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{K}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{K}$, $\boldsymbol{\mathit{V}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{V}$, $\boldsymbol{\mathit{Q}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{K}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{K}$, $\boldsymbol{\mathit{V}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{V}$. Then, we apply the forward pass of the efficient parallel training algorithm (Algorithm 1) to get $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$. For the UQ-KV cache, since we already have $\boldsymbol{\mathit{Q}}^{U}_{L}$, $\boldsymbol{\mathit{K}}^{C}_{L}$ and $\boldsymbol{\mathit{V}}^{C}_{L}$, we only need to obtain $\boldsymbol{\mathit{U}}^{L}$. This can be done similarly to FlashAttention-2 [5] due to the similarity between (2) and standard causal attention. The complete prefilling procedure is given in Algorithm 4.
The analysis above leads to the following theorem.
@@
Given inputs $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$, prefilling has computational complexity $O(Ldd_{\rm model}+L^{2}d)$ and space complexity $O(Ld)$. During the decoding stage, when generating the $t$-th token, the computational complexity is $O(td+dd_{\rm model})$ and the UQ-KV cache requires $O(td)$ memory.
@@ 2.5 Multi-Head Causal Attention with Lookahead Keys
where $\boldsymbol{\mathit{W}}^{O}\in\mathbb{R}^{nd\times d_{\rm model}}$ is a learnable matrix. The formula for forward pass in parallel training and more details of multi-head CASTLE are in Appendix 6.
@@ 3.1 Experimental Setup
Our baseline follows the improved Transformer architecture with SwiGLU [20], Rotary Positional Embeddings [22], and RMSNorm [29] following LLaMA [24]. We train models on four scales from scratch: small (0.16B), medium (0.35B), large (0.75B) and XL (1.3B). The medium, large and XL baseline models mirror the configuration of GPT-3 [3]. For the small setting, we increase the number of heads and the hidden dimension relative to the original GPT-3 configuration to better align parameter counts between standard causal attention and CASTLE. To isolate the effect of the attention mechanism, we replace standard causal attention with dynamic attention with evolving keys while keeping all other components unchanged for a fair comparison. We use the AdamW optimizer [14] and follow the training recipe of [6]. All models are trained on FineWeb-Edu dataset [15] for 50B tokens. Further details of experimental setup can be found in Appendix 9.
@@ 3.2 Training & Validation Loss
We report training and validation loss and perplexity in Table 1. Training loss and validation loss curves are shown in Figure 5, Figure 7, Figure 8 and Figure 9. CASTLE consistently outperforms the baseline models in both training and validation loss across all model scales.
Specifically, after training 50B tokens, CASTLE outperforms baselines across all model scales and achieves validation losses that are 0.0059, 0.0245, 0.0356, and 0.0348 lower than the baseline for the small, medium, large, and XL models, respectively. The performance gains are particularly significant in the medium, large, and XL models.
@@
Furthermore, as shown in Table 4, CASTLE has the same or fewer parameters than their baseline counterparts. This further underscores CASTLE’s effectiveness in improving model performance.
Table 1: Training and validation loss and perplexity for models with standard causal attention and CASTLE. We use “S”, “M”, “L”, “XL” to denote model scales. Each model is trained on FineWeb-Edu for 50B tokens. Significant improvements in loss or perplexity are in bold.
@@
Loss PPL Loss PPL
Baseline-S 2.795 16.364 2.798 16.411
CASTLE-S 2.789 16.259 2.792 16.315
Baseline-M 2.641 14.030 2.639 14.004
CASTLE-M $\bm{2.616}$ $\bm{13.684}$ $\bm{2.615}$ $\bm{13.665}$
Baseline-L 2.513 12.346 2.507 12.269
CASTLE-L $\bm{2.476}$ $\bm{11.897}$ $\bm{2.472}$ $\bm{11.840}$
Baseline-XL 2.430 11.360 2.426 11.309
CASTLE-XL $\bm{2.401}$ $\bm{11.031}$ $\bm{2.391}$ $\bm{10.922}$
Figure 5: Training and validation loss curves of XL models. Training loss curve is smoothened with a moving window of 2000 training steps. Validation loss is evaluated every 100 training steps on 40M tokens, and its curve is smoothened by a moving window of 20 evaluation intervals. Loss curves for the small, medium and large models can be found in Figure 7, Figure 8 and Figure 9 of Appendix 9.3. After 50B training tokens, CASTLE-XL achieves a $\bm{0.0294}$ lower training loss and a $\bm{0.0348}$ lower validation loss compared to Baseline-XL.
@@ 3.3 Downstream Evaluation
We evaluate CASTLE on a diverse set of downstream benchmarks, including ARC [27], BoolQ [4], HellaSwag [28], MMLU [10], OBQA [16], PIQA [2], Winograde [19] using lm-evaluation-harness [9]. Table 1 reports normalized accuracy for ARC-Challenge, HellaSwag, OBQA, and PIQA, and standard accuracy for the other benchmarks. Results are reported in Table 2 (0-shot) and Table 3 (5-shot). Across all model scales and evaluation settings, CASTLE consistently outperforms the baseline in average downstream accuracy. In the 0-shot setting, CASTLE demonstrates robust improvements on ARC-E, BoolQ, HellaSwag, and OBQA across all scales, highlighting enhanced reasoning and commonsense capabilities. In the 5-shot setting, CASTLE further strengthens performance, with consistent gains on ARC-C across all scales, underscoring its strong few-shot generalization on reasoning and knowledge integration abilities. These findings accompany lower loss and perplexity in Section 3.2 demonstrate that CASTLE not only lowers perplexity but also translates these gains into stronger performance on diverse downstream tasks.
@@
CASTLE-S 26.19 $\bm{56.69}$ $\bm{59.85}$ $\bm{36.28}$ $\bm{23.00}$ $\bm{31.60}$ $\bm{64.25}$ 52.25 $\bm{43.76}$
Baseline-M 28.58 60.90 53.61 43.01 23.21 33.40 $\bm{67.95}$ 50.91 45.20
CASTLE-M $\bm{30.20}$ $\bm{61.36}$ $\bm{58.01}$ $\bm{43.24}$ $\bm{25.34}$ $\bm{34.60}$ $\bm{67.95}$ $\bm{52.64}$ $\bm{46.67}$
Baseline-L $\bm{32.59}$ 65.07 57.49 47.45 23.57 32.60 70.51 50.75 47.50
CASTLE-L 32.34 $\bm{65.15}$ $\bm{57.65}$ $\bm{47.87}$ $\bm{24.51}$ $\bm{35.60}$ $\bm{70.78}$ $\bm{53.51}$ $\bm{48.43}$
@@
Baseline-S 25.68 $\bm{54.97}$ 56.09 33.81 $\bm{25.54}$ 28.20 63.98 $\bm{52.57}$ 42.60
CASTLE-S $\bm{26.02}$ 54.25 $\bm{57.13}$ $\bm{35.24}$ 25.22 $\bm{29.80}$ $\bm{64.53}$ 50.99 $\bm{42.90}$
Baseline-M 31.06 62.46 48.47 42.83 $\bm{25.22}$ 33.00 68.39 51.78 45.40
CASTLE-M $\bm{32.17}$ $\bm{64.06}$ $\bm{54.62}$ $\bm{43.47}$ $\bm{25.22}$ $\bm{33.80}$ $\bm{69.48}$ $\bm{52.49}$ $\bm{46.91}$
Baseline-L 33.36 63.64 $\bm{59.24}$ 46.16 $\bm{26.82}$ 33.40 69.53 $\bm{54.06}$ 48.28
CASTLE-L $\bm{37.37}$ $\bm{67.89}$ 50.95 $\bm{47.71}$ 26.11 $\bm{34.20}$ $\bm{70.18}$ $\bm{54.06}$ $\bm{48.56}$
Baseline-XL 35.58 65.78 61.07 50.84 $\bm{26.71}$ 36.20 $\bm{71.27}$ 52.72 50.02
CASTLE-XL $\bm{39.08}$ $\bm{70.24}$ $\bm{62.60}$ $\bm{51.63}$ 24.16 $\bm{37.40}$ 71.00 $\bm{58.33}$ $\bm{51.80}$

### v2

@@ 2.4 Efficient Inference with UQ-KV Cache
In this section, we introduce the inference algorithm which has unified forms for CASTLE and CASTLE-SWL. We first introduce the decoding algorithm. The decoding algorithm consists of the following 2 steps: updating step and combining step. Fix $t\in\left\{1,2,\dots,L\right\}$, and consider generating the $(t+1)$-th token.
@@
where $\boldsymbol{\mathit{M}}^{U}_{t}]_{:,t}$ is the $t$-th column of $\boldsymbol{\mathit{M}}^{U}_{t}$. The proof of (13) is given in Appendix 9. Next, we discuss the caching strategy and FLOPs of the updating step.
@@
For the prefilling stage, let the prompt length be $L$ and inputs $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$. We first compute $\boldsymbol{\mathit{Q}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{K}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{K}$, $\boldsymbol{\mathit{V}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{V}$, $\boldsymbol{\mathit{Q}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{K}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{K}$, $\boldsymbol{\mathit{V}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{V}$. Then, we apply the forward pass of the efficient parallel training algorithm (Algorithm 1) to get ${\rm Attention}(\boldsymbol{\mathit{X}}^{L})$. For the UQ-KV cache, since we already have $\boldsymbol{\mathit{Q}}^{U}_{L}$, $\boldsymbol{\mathit{K}}^{C}_{L}$ and $\boldsymbol{\mathit{V}}^{C}_{L}$, we only need to obtain $\boldsymbol{\mathit{U}}^{L}$. This can be done similarly to FlashAttention-2 [5] due to the similarity between (3) and standard causal attention. The complete prefilling procedure is given in Algorithm 4. The analysis above leads to the following theorem.
@@
Given inputs $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$, for both CASTLE and CASTLE-SWL, prefilling has computational complexity $O(Ldd_{\rm model}+L^{2}d)$ and space complexity $O(Ld)$, and during the decoding stage, when generating the $t$-th token, the computational complexity is $O(td+dd_{\rm model})$ and the UQ-KV cache requires $O(td)$ memory.
@@ 2.5 Multi-Head Causal Attention with Lookahead Keys
where $\boldsymbol{\mathit{W}}^{O}\in\mathbb{R}^{nd\times d_{\rm model}}$ is a learnable matrix. The formula for forward pass in parallel training and more details of multi-head CASTLE are in Appendix 7. Multi-head CASTLE-SWL shares the same formulation and parameter count as CASTLE, and is omitted here to avoid redundancy.
@@ 3.1 Experimental Setup
We use the nanoGPT [13] code base. Our baseline follows the improved Transformer architecture with SwiGLU [22], Rotary Positional Embeddings [24], and RMSNorm [32] following LLaMA [26]. We train models on four scales from scratch: small (0.16B), medium (0.35B), large (0.75B) and XL (1.3B). The medium, large and XL baseline models mirror the configuration of GPT-3 [3]. For the small setting, we increase the number of heads and the hidden dimension relative to the original GPT-3 configuration to better align parameter counts between standard causal attention and CASTLE. CASTLE-SWL shares identical configurations with CASTLE of each model scale. To isolate the effect of the attention mechanism, we replace standard causal attention with CASTLE or CASTLE-SWL while keeping all other components unchanged for a fair comparison. We use the AdamW optimizer [16] and follow the training recipe of [6]. All models are trained on FineWeb-Edu dataset [17] for 50B tokens. Further details of experimental setup can be found in Appendix 5.1.
@@ 3.2 Training & Validation Loss
We report training and validation loss and perplexity in Table 1. Training loss and validation loss curves are shown in Figure 6, Figure 7, Figure 8 and Figure 9. CASTLE and CASTLE-SWL consistently outperform the baselines in both training and validation loss across all model scales.
Specifically, after training 50B tokens, CASTLE outperforms baselines across all model scales and achieves validation losses that are 0.0059, 0.0245, 0.0356, and 0.0348 lower than the baseline for the small, medium, large, and XL models, respectively.
@@
CASTLE-SWL matches CASTLE’s performance and its validation loss outperforms baselines by 0.0084, 0.0241, 0.0366, 0.0369, respectively. The performance gains are particularly significant in the medium, large, and XL models.
Furthermore, as shown in Table 4, both CASTLE and CASTLE-SWL have the same or fewer parameters than their baseline counterparts. This further underscores CASTLE and CASTLE-SWL’s effectiveness in improving model performance.
Table 1: Training and validation loss and perplexity for models with CASTLE, CASTLE-SWL and standard causal attention. We use “S”, “M”, “L”, “XL” to denote model scales. Each model is trained on FineWeb-Edu for 50B tokens. The best and second best results (when showing clear improvements upon baselines) are highlighted in bold and underline, respectively.
@@
$n_{\text{params}}$ Loss PPL Loss PPL
Baseline-S 160M 2.795 16.364 2.798 16.411
CASTLE-S 160M 2.789 16.259 2.792 16.315
CASTLE-SWL-S 160M 2.786 16.213 2.790 16.273
Baseline-M 353M 2.641 14.030 2.639 14.004
CASTLE-M 351M $\bm{2.616}$ $\bm{13.684}$ $\bm{2.615}$ $\bm{13.665}$
CASTLE-SWL-M 351M 2.618 13.708 2.615 13.670
Baseline-L 756M 2.513 12.346 2.507 12.269
CASTLE-L 753M 2.476 11.897 2.472 11.840
CASTLE-SWL-L 753M $\bm{2.476}$ $\bm{11.890}$ $\bm{2.471}$ $\bm{11.832}$
Baseline-XL 1.310B 2.430 11.360 2.426 11.309
CASTLE-XL 1.304B $\bm{2.401}$ $\bm{11.031}$ 2.391 10.922
CASTLE-SWL-XL 1.304B 2.401 11.036 $\bm{2.389}$ $\bm{10.900}$
Figure 6: Training and validation loss curves of XL models. Training loss curve is smoothened with a moving window of 2000 training steps. Validation loss is evaluated every 100 training steps on 40M tokens, and its curve is smoothened by a moving window of 20 evaluation intervals. Loss curves for the small, medium and large models can be found in Figure 7, Figure 8 and Figure 9 of Appendix 5.3. After 50B training tokens, CASTLE-XL achieves a 0.0294 lower training loss and a 0.0348 lower validation loss compared to Baseline-XL, while CASTLE-SWL-XL achieves a 0.0290 lower training loss and a 0.0369 lower validation loss compared to Baseline-XL
@@ 3.3 Downstream Evaluation
We evaluate CASTLE and CASTLE-SWL on a diverse set of downstream benchmarks, including ARC [29], BoolQ [4], HellaSwag [31], MMLU [11], OBQA [18], PIQA [2], Winograde [21] using lm-evaluation-harness [10]. We report normalized accuracy for ARC-Challenge, HellaSwag, OBQA, and PIQA, and standard accuracy for the other benchmarks. Results are reported in Table 2 (0-shot) and Table 3 (5-shot). Across all model scales and evaluation settings, both CASTLE and CASTLE-SWL consistently outperform the baselines in average downstream accuracy under both 0-shot and 5-shot settings. These findings accompany lower loss and perplexity in Section 3.2 demonstrate that CASTLE and CASTLE-SWL not only lower perplexity but also translate these gains into stronger performance on diverse downstream tasks.
@@
CASTLE-S 26.19 $\bm{56.69}$ $\bm{59.85}$ $\bm{36.28}$ $\bm{23.00}$ 31.60 64.25 52.25 $\bm{43.76}$
CASTLE-SWL-S 26.62 54.12 59.60 35.97 22.87 $\bm{32.60}$ $\bm{64.31}$ 50.51 43.33
Baseline-M 28.58 60.90 53.61 43.01 23.21 33.40 67.95 50.91 45.20
CASTLE-M $\bm{30.20}$ 61.36 $\bm{58.01}$ 43.24 $\bm{25.34}$ $\bm{34.60}$ 67.95 52.64 $\bm{46.67}$
CASTLE-SWL-M 29.52 $\bm{61.41}$ 57.03 $\bm{43.76}$ 23.29 33.00 $\bm{68.12}$ $\bm{53.83}$ 46.24
Baseline-L 32.59 65.07 57.49 47.45 23.57 32.60 70.51 50.75 47.50
CASTLE-L 32.34 $\bm{65.15}$ 57.65 47.87 $\bm{24.51}$ 35.60 $\bm{70.78}$ 53.51 48.43
CASTLE-SWL-L $\bm{32.76}$ 63.51 $\bm{60.95}$ $\bm{48.50}$ 23.72 $\bm{36.00}$ 70.02 $\bm{54.85}$ $\bm{48.79}$
@@
Baseline-S 25.68 54.97 56.09 33.81 $\bm{25.54}$ 28.20 63.98 $\bm{52.57}$ 42.60
CASTLE-S 26.02 54.25 $\bm{57.13}$ 35.24 25.22 29.80 64.53 50.99 42.90
CASTLE-SWL-S $\bm{27.39}$ $\bm{56.19}$ 56.09 $\bm{35.46}$ 25.34 $\bm{30.20}$ $\bm{64.85}$ 51.22 $\bm{43.34}$
Baseline-M 31.06 62.46 48.47 42.83 25.22 33.00 68.39 51.78 45.40
CASTLE-M 32.17 $\bm{64.06}$ $\bm{54.62}$ 43.47 25.22 $\bm{33.80}$ $\bm{69.48}$ 52.49 46.91
CASTLE-SWL-M $\bm{32.85}$ 63.85 54.19 $\bm{44.18}$ $\bm{26.43}$ $\bm{33.80}$ 69.04 $\bm{53.43}$ $\bm{47.22}$
Baseline-L 33.36 63.64 59.24 46.16 $\bm{26.82}$ 33.40 69.53 $\bm{54.06}$ 48.28
CASTLE-L $\bm{37.37}$ $\bm{67.89}$ 50.95 47.71 26.11 34.20 $\bm{70.18}$ $\bm{54.06}$ 48.56
CASTLE-SWL-L 36.26 65.53 $\bm{60.58}$ $\bm{48.55}$ 24.70 $\bm{34.80}$ 69.10 53.67 $\bm{49.15}$
Baseline-XL 35.58 65.78 61.07 50.84 $\bm{26.71}$ 36.20 71.27 52.72 50.02
CASTLE-XL $\bm{39.08}$ $\bm{70.24}$ $\bm{62.60}$ 51.63 24.16 $\bm{37.40}$ 71.00 $\bm{58.33}$ 51.80
CASTLE-SWL-XL 38.99 70.08 61.74 $\bm{52.35}$ 25.85 37.20 $\bm{72.52}$ 56.75 $\bm{51.93}$